AS/AD Model.
AD: P = 90 - 3Y.
AS: P = 15.
PO: Y = 30.
What is the Inflation?
1
Expert's answer
2016-05-13T12:37:03-0400
If in AS/AD Model AD function is P = 90 - 3Y, AS function is P = 15, potential output (PO) is Y = 30, then in equilibrium:
AS = AD, so 90 - 3Y = 15, Y = 25.
If at PO Y = 30, then there is a recessionary gap (real GDP is lower than potential GDP), so there is no inflation in this case.